OpenStudy (anonymous):

Find the standard form of the equation of the parabola with a focus at (0, -10) and a directrix at y = 10. (5 points) y = -1/10x2 y =-1/40 x2 y2 = -40x y2 = -10x

OpenStudy (anonymous):

did you make a sketch?

OpenStudy (anonymous):

no... @sourwing

OpenStudy (anonymous):

perhaps you should make a sketch now?

OpenStudy (anonymous):

how??

OpenStudy (anonymous):

@sourwing

OpenStudy (anonymous):

there is a blue [draw] button below. Click that and draw

OpenStudy (anonymous):

i don't know how to draw.

OpenStudy (anonymous):

@sourwing

OpenStudy (michele_laino):

the general equation for a parabole whose focus is: F=(0, p/2) and whose directrix is: y= -p/2, is: \[\Large {x^2} = 2py\] now, using your data we have: \[\Large \frac{p}{2} = - 10\] so, what is p=...? then substitute that value of p, into my equation above, so you will get your answer

OpenStudy (michele_laino):

parabola*

OpenStudy (anonymous):

-5 !!!! @Michele_Laino